OpenStudy (anonymous):

A 0.20 kg block of a substance requires 3.59 kJ of heat to raise its temperature by 20 K. What is the substance? *will draw table* iron, aluminum, gold, or copper?

OpenStudy (michele_laino):

we have to compute the specific heat of that substance, and then compare it with the values of your table

OpenStudy (michele_laino):

\[c = \frac{1}{m}\frac{{\Delta Q}}{{\Delta t}} = \frac{1}{{0.2}}\frac{{3.59 \times {{10}^3}}}{{20}}\]

OpenStudy (anonymous):